# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v1

@@ 5 Evaluation
  • Microbenchmarks: We conduct experiments to show the
@@
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study how sensitive is our algorithm to different values of parameter $p$.
5.1 RQ1: Using PRI To Choose Epsilon
@@
To determine whether the PRI help the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. Our goal is to show that by presenting the PRI to the participants, they will be more consistent on the $\epsilon$ they choose compared to if they don’t have the PRI. To this end, we conducted a between-subject user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey, because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
5.1.1 Recruitment and data collection
@@
We recruited 100 participants using Prolific[1]. To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP Address, location data, and contact information are not being recorded.
@@
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains DP is generally achieved by adding random noise to the query output, and $\epsilon$ is inversely proportional to the amount of noise added.
@@
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Out of the 100 participants we recruited, 56 of them passed the quiz and their responses are used for the study analysis. This quiz ensures that participants have understood the principles of DP and thus ensures the validity of the user study.
5.1.2 Study procedure
@@
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group and 27 participants are in the treatment group. The survey presents a sample from the UCI Adult dataset[2] and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order and asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 and each interval being 0.5. The participant’s task is to choose an $\epsilon$ value that offers equal privacy protection to all people in the dataset while preserving as much accuracy as possible.
@@
For the control group, the participant is presented with a table that shows the range of possible output for each $\epsilon$, indicating output accuracy, along with the ground truth (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
@@
For the treatment group, the participant is presented with the same table, but with an additional column that includes a plot showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI). Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. We explain beforehand what the PRI is at a high level, and that a smaller range represents more similar PRIs among people in the dataset. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
5.1.3 Results
Figure 5: $\epsilon$ chosen by each participant
@@ 5.1.3 Results
We compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups. Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control and treatment group for each query. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to offer equal privacy protection while ensuring high accuracy, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to pick.
@@
In addition, we ran the Mann–Whitney test [37] on the $\epsilon$ values in the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
In the follow-up question for the treatment group, all 27 participants answered yes when asked whether they used the PRIs when choosing $\epsilon$, confirming our hypothesis that presenting the PRIs helped participants choose the $\epsilon$ that satisfies the task goal.
@@
Figure 6: $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
@@
Does knowing DP matter? Another interesting question is whether the participant’s a priori familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they don’t know anything about DP and 29 participants answered they know about DP.
@@
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$, instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs - by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
Figure 7: $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, 8 participants chose larger $\epsilon$ (3.5, 4, and 4.5). Although we cannot be certain since the responses are anonymous, one possible explanation might be that for the large $\epsilon$, the majority of PRI are already small and similar to each other. Only a small portion of PRI have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
@@
5.1.4 Threats to validity
External validity. We recruited participants from diverse backgrounds across the world via Prolific, with few prescreen criteria such that the participants can complete the study without difficulty.
Internal validity. To ensure all participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. We only allowed the participant to take the quiz once to minimize the chance of participants abusing it if they got multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
5.2 RQ2: DP On Escrow-based Platform

### v2

@@ 3.4. End-to-End DP Guarantee
The privacy odometer gives the same privacy bound as sequential composition; since the algorithm uses $\epsilon$-DP mechanism, $\delta_{g}$ (the global failure parameter in $(\epsilon,\delta)$-differential privacy) can be set to 0. The privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}$ after the algorithm answers each query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $\epsilon_{c}$-differential privacy.
It is worth noting that although  (Rogers et al., 2016) showed the privacy odometer gives a worse bound on privacy loss than the standard advanced composition theorem (Dwork et al., 2010), more recently (Whitehouse et al., 2023) proved that in the fully-adaptive setting the privacy odometer can achieve the same asymptotic bound as advanced composition under certain conditions. For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition.
3.5. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel. Thus, we implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 6.3.1 that the algorithm is both efficient and scalable.
Expressing privacy preferences. By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$. A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records. For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection. This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The PRI complements the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4. Helping Analysts Interpret Utility
In this section, we concentrate on solutions to help analysts interpret the utility of DP output. We revisit Challenge 2 (Section 4.1) and present a solution we call utility signaling protocol (Section 4.2).
4.1. Challenge Overview and Key Insight
Controllers work on a platform that has access to the individuals’ data and that is responsible for answering analysts’ queries using DP. An analyst submits a set of queries $Q=\{q\mid q\in\mathcal{Q}\}$ ($\mathcal{Q}$ is the query space) to the platform, and (if a suitable $\epsilon$ is found for each query) receives a set of noisy query output $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n},q\in Q\}$ ($\mathcal{M}$ is a DP mechanism and $\mathcal{X}$ is the data universe). A well-known challenge of using DP is that analysts do not know the utility of $S$ (Hotz et al., 2022). Concretely, analysts do not know the effect of using $S$ on their downstream task, which is the purpose for analysts to query the data in the first place.
Example 1/3. An analyst queries census data at the county level. The set of query output $S$ is aggregate data on population count, education level, and housing prices that are released using DP. The analyst uses $S$ to train a binary classifier to predict the income level of each county and obtains an accuracy of 68%. The analyst does not know if the (low) accuracy is due to the effect of DP or inappropriate training data.
The difficulty in interpreting the utility of $S$ makes it challenging for analysts to understand how the DP output will impact their downstream tasks; this is one of the core issues surrounding the discussion of the use of DP for the US Census (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
Observation. Analysts use the DP output $S$ for some downstream task, such as a machine learning or inference task. Downstream tasks often have an associated evaluation function, $g(\cdot)\rightarrow\mathbb{R}$, which returns a number indicating the quality of the downstream task when evaluated on a certain input. In the example above, the downstream task is to train a binary classifier, the evaluation function measures the quality of the classifier trained on DP output and returns 68% as the quality indication (accuracy in this case).
Insight: signaling DP effect. Analysts would learn the impact of DP if they had access to both the result of the evaluation function on the DP output, $g(S)$ [^2] We slightly abuse notation for simplification. and the raw output, $g(S^{\prime})$, where $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n},q\in Q\}$ (the set of query output without using DP).
Example 2/3. Besides having access to the quality of the classifier on the DP census aggregates (68%), the analyst also gains access to the quality of the classifier if it had been trained on the raw census aggregates $S^{\prime}$, say 80%. By comparing 68% ($g(S)$) and 80% ($g(S^{\prime})$), the analyst learns the effect of added DP noise on their downstream task.
Gaining access to $g(S)$ and $g(S^{\prime})$ helps analysts interpret the impact of DP on their downstream tasks. The remaining challenge is how to provide $g(S^{\prime})$ to the analysts without giving them access to $S^{\prime}$, which is the raw query output and cannot be released to the analyst without violating the privacy of individuals in the dataset.
4.2. The utility signaling protocol
Rather than the platform directly sending $S^{\prime}$ to the analyst, which would break individuals’ privacy, the analyst submits their downstream task and associated evaluation function $g(\cdot)$ to the platform. Upon successful inspection by the controller, the platform executes the downstream task and evaluation function on behalf of the analyst, producing $g(S)$ and $g(S^{\prime})$. Last, if the controller approves, the platform releases both $g(S)$ and $g(S^{\prime})$ to the analyst, thus signaling the impact of DP on the analysts’ task without revealing $S^{\prime}$.
Example 3/3. Instead of only submitting the queries to the platform, with the utility signaling protocol, the analyst submits their queries, the downstream task, and the evaluation function. The controller inspects each and allows the platform to signal the impact of using DP to the analyst by releasing both utility results 68% and 80%.
4.2.1. Protocol Step by Step
Figure 4. End-to-end example of utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
Figure 4 illustrates the utility signal protocol, which we describe next:
0. Platform. The platform has access to the dataset $x$ of individuals. It also has the capacity to run queries using DP. It may optionally implement the Find-$\epsilon$-from-PRI algorithm presented in Section 3.2333While we envision the algorithm being deployed to help controllers, both contributions are orthogonal.
1. Analyst submits queries to the platform. The platform notifies the controller.
2. Controller inspects queries. The controller inspects the analyst’s queries and allows the platform to execute the queries as long as a suitable $\epsilon$ is selected. The platform withholds both the set of DP output $S$ and raw output $S^{\prime}$ without releasing anything to the analyst by default.
3. Platform releases $S$ upon controller’s approval. The platform notifies the controller after computing $S$ and $S^{\prime}$. The controller gives permission for the platform to release $S$ to the analyst.
Up to this point, the protocol provides the analyst with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help the analyst understand the impact of DP on their downstream task.
4. Analyst submits downstream task and evaluation function $g(\cdot)$ to the platform. The platform notifies the controller.
5. Controller inspects downstream task and evaluation function. The controller inspects the implementation of the analyst’s downstream task and evaluation function $g(\cdot)$, and decides whether it is appropriate to release $g(S)$ and $g(S^{\prime})$. With the controller’s approval, the platform executes the downstream task and evaluation function $g(\cdot)$ on both $S$ and $S^{\prime}$, withholding the results by default.
6. Platform releases $g(S)$ and $g(S^{\prime})$ upon controller’s approval. The controller can inspect $g(S^{\prime})$ one last time to ensure it is appropriate to release; $g(S)$ preserves DP since it is a post-processing (Dwork et al., 2014) step of DP output $S$. If approved, the platform releases $g(S)$ and $g(S^{\prime})$ to the analyst, who learns the impact of DP on their downstream task without ever gaining access to $S^{\prime}$ directly.
4.2.2. Privacy implications of releasing $g(S^{\prime})$
Any data release leaks information. When releasing an output that is subject to DP, such as $g(S)$, controllers understand the consequences on the individual’s privacy. Conversely, when releasing output not subject to DP, such as $g(S^{\prime})$, controllers must assess whether doing so maintains individuals’ privacy. That ultimately depends on the dataset $x$, the output $S^{\prime}$, the implementation of the downstream task and evaluation function, $g(\cdot)$, and the adversary model. The scope of the utility signaling protocol is to enable analysts to understand the impact of using DP, but not on understanding, controlling, and determining when it is safe to release $g(S^{\prime})$, which is ultimately at the discretion of the controllers and an interesting avenue of future work.
5. Escrow-Based DP Platform
We curate a set of requirements for a DP platform that implements the Find-$\epsilon$-from-PRI algorithm and the utility signaling protocol in Section 5.1. In Section 5.2, we describe the implementation of a platform built on top of a data escrow that meets these requirements.
5.1. Platform Requirements
The solutions to Challenge 1 (the Find-$\epsilon$-from-PRI algorithm), to Challenge 2 (the utility signaling protocol), and the nature of processing over private data call for a platform that meets the following requirements:
  • R1. Control dataflows: Given the three relevant entities (analyst, platform, and controller), a dataflow indicates data movement between any two of those entities. A requirement for the platform is to ensure that only permitted dataflows take place. Concretely, controllers indicate what dataflows take place, and the platform enforces the controller’s decision. In the context of our work, analysts can access for example the DP output $S$ after the controller approves, and they cannot access PRIs or any other intermediates not explicitly approved by the controller.
  • R2. Secure platform: Adversarial agents may break the requirement of controlling dataflows by breaking into the platform and forcing unintended dataflows. This is an assumed risk in the central DP model, which can be reduced by using a local DP deployment (Near et al., 2021). Unfortunately, local DP reduces the accuracy of the output compared to central DP. A requirement for the platform is to ameliorate the risk of unintended dataflows.
  • R3. Protect against adversarial downstream and evaluation functions: Adversarial analysts could submit downstream tasks and evaluation functions crafted to release more information than needed. While controllers must inspect these functions before agreeing to release the output and they are the ultimate arbiter, a requirement of the platform is to facilitate the controllers’ job.
  • R4. High Performance: The Find-$\epsilon$-from-PRI algorithm increases processing overhead. Simultaneously, security safeguards to address R2 are often at odds with high-performance computation. Together, these two facts risk making the platform impractical. A requirement is to implement the solutions to Challenge 1, Challenge 2, and meet the requirements R1, R2, and R3 while maintaining low overhead.
We now explain the implementation of a platform that addresses these three requirements.
5.2. Platform Design and Implementation
To demonstrate the feasibility of the solutions to Challenge 1 and Challenge 2, we describe how to implement the solutions above on a platform that leverages a data escrow (Xia et al., 2022) architecture.
Data escrow overview. A data escrow is a system that acts as an intermediary for storing, processing, and releasing data. Different entities interact within the escrow. Controllers register data and control the dataflows. The data is never stored or computed in the clear using end-to-end encryption. Analysts run computations on registered data in the escrow. For any data or its derived data product to leave the escrow, controllers have to create a policy on the escrow. The policies are tamper-proof since the escrow validates the identity of the controller who creates the policy and the policy remains encrypted, thus ensuring the controller fully controls the dataflow.
Meeting R1. No raw data (e.g., $x$) or derived data product (e.g., PRIs, $g(S)$, $g(S^{\prime})$) leave the escrow without an explicit policy that permits that. Policies then determine what dataflows take place. Controllers are in charge of writing policies.
Meeting R2. Data is encrypted end-to-end, even during computation. Adversaries that break into the platform cannot read plain-text data. In other words, no unintended dataflows take place.
Meeting R3. The platform can let analysts submit their downstream task and associated evaluation function, and put the burden on the controller to verify the implementation of those functions. Or it can offer standard implementation of downstream tasks and evaluation functions and let analysts call these, thus requiring scrutinizing the functions once and using them many times. The data escrow permits the deployment of programs expressing a desired functionality. In specific DP scenarios, e.g., the US Census, it is conceivable that one could provide programs implementing (at least) the more standard functionalities and query interfaces, and rely on controllers to scrutinize only one-off tasks.
Meeting R4. End-to-end data encryption (necessary to meet R2) can be achieved in many ways; for instance, DP platforms leveraging Secure Multiparty Computation (Narayan and Haeberlen, 2012; Bater et al., 2016, 2018; Wang et al., 2022; Roth et al., 2019), Homomorphic Encryption (Agarwal et al., 2018; Roy Chowdhury et al., 2020), Oblivious RAM (Bogatov et al., 2021) and Trusted Execution Environment (Nguyen et al., 2020). Our evaluation prototype uses AMD Secure Encrypted Virtualization (Kaplan et al., 2016) because it permits running unmodified software while maintaining the integrity and confidentiality of memory contents. The escrow design protects access to external resources (e.g., network and disk), ensuring data is encrypted transparently to any application. As we show in the evaluation (Section 6.2), this solution virtually removes the processing overhead. In addition, we implement a series of optimizations to the Find-$\epsilon$-from-PRI algorithm to demonstrate its efficiency and scalability when run in the escrow (Section 6.3.1).
6. Evaluation
Our evaluation shows our solutions address Challenge 1 to help controllers, and Challenge 2 and 3 using an escrow-based platform. We answer the following research questions:
  • RQ1: Does presenting the PRI help data controllers choose $\epsilon$?
[…]
  • Microbenchmarks: We conduct experiments to show that
@@ 6. Evaluation
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study the effect of varying parameter $p$ in the algorithm.
6.1. RQ1: Using PRI To Choose Epsilon
@@
To determine whether PRI helps the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. To this end, we conducted a between-subjects user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
6.1.1. Recruitment and data collection
@@
We recruited 100 participants using Prolific (pro, [n.d.]). To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP address, location data, and contact information are not being recorded.
@@
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains that DP is generally achieved by adding random noise to the query output, and larger $\epsilon$ means less noise will be added and vice versa.
@@
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Of the 100 participants we recruited, 56 passed the quiz and their responses were used for the study analysis. This quiz ensures that participants have understood the principles of DP, thus strengthening the validity of the user study.
6.1.2. Study procedure
@@
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group. The survey presents a sample from the UCI Adult dataset (uci, [n.d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order:
  • The number of people with income greater than 50K and are less than 40 years old.
  • The number of people who have a Master’s degree or above and have income $\leq$ 50k, grouped by each race.
For each query, the survey asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 in an interval of 0.5. Since we want all participants to have a clear task goal in mind, the survey includes a description of the task: choose $\epsilon$ to protect individuals’ privacy while yielding as accurate answers as possible.
@@
For the control group, the participant is presented with a table showing the range of possible output for each $\epsilon$ to indicate output accuracy and the raw (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
@@
For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI. Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
6.1.3. Results
Figure 5. $\epsilon$ chosen by each participant
If the intervention (showing the PRIs) is effective, we expect this would manifest as more consistent answers among participants, since all participants were given the same training and the same task. Therefore, we compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups.
@@ 6.1.3. Results
Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control or treatment group. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to protect individuals and release useful output, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to choose.
@@
We ran Mann–Whitney test  (McKnight and Najab, 2010) on the $\epsilon$ values chosen from the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
Last, we asked participants in the treatment group if they used the PRIs to complete the task: all 27 participants responded they did use that information, confirming they were indeed “treated”.
@@
Figure 6. $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
@@
Does knowing DP matter? Another interesting question is whether the participant’s prior familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they do not know anything about DP and 29 participants answered they know about DP.
@@
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
Figure 7. $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, eight participants chose larger $\epsilon$ (3.5, 4, and 4.5). Since the responses are anonymous, we cannot directly ask participants about their decision-making after the survey ends. We deduce one possible explanation might be that for the large $\epsilon$, the majority of PRIs are already small and similar, and only a small portion of PRIs have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
@@
6.1.4. Threats to validity
External validity: We recruited participants from diverse backgrounds worldwide via Prolific, with few prescreen criteria so that the participants could complete the study without difficulty. Internal validity: To ensure participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. Participants can only take the quiz once to minimize the chance of them abusing it if they have multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
6.2. RQ2: DP On Escrow-based Platform